OpenStudy (anonymous):

Which statement shows the distributive property? A. 13 · (k + 10) = 13 · k + 13 · 10 B. 13 · (k + 10) = 13 · k + 10 C. 13 · (k + 10) = 13 · k · 10 D. 13 · (k + 10) = (k + 10) · 13

OpenStudy (anonymous):

a(b+c) = a(b) + a(c) which choice shows that?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

2(x - 6) = 2(x) - 2(6)